all 87 comments

sorted by: hot top controversial new old
[–] 59 points 2 months ago (1 child)

Help, my mind is caught in a paradoxical loop!

  • source
  • hideshow 1 child comment
  • [–] 34 points 2 months ago (3 children)

    Going with D. It'd be 25% for a normal question, but since it's the second 25% this provides some evidence that I read all the answers.

  • source
  • hideshow 3 child comments
  • [–] 13 points 2 months ago (2 children)

    2 out of 4 answers are the same, and if we say that 25% is correct then you actually have a much higher chance of selecting the correct one. I’m sure there’s some statistics crap to say it’s not 50%, but if half your random choices are correct then that seems right to me. So now 50% is correct, but it’s only 1 in 4 and the odds go back to 25%. So now it’s 0%, except that that’s ALSO on there so now the correct answer is 25% again. The problem is that ever single time 25% becomes the correct answer you now have a 50% chance to get it right if chosen randomly and it’s wrong all over again.

  • source
  • parent
  • hideshow 2 child comments
  • [–] 5 points 2 months ago* (1 child)

    Yep, that's definitely the thought process the question-asker intended. I'm sidestepping that slightly: if an answer is chosen completely at random then one should be able to completely ignore the text of the question and answer, and roll a four-sided die, leading to 25%. I agree that strategy assumes that there's a single correct answer. And the typical expectation of a multiple choice question is to only have a single correct answer. The question itself says "the answer" not "answer(s)", that's perhaps a partial support of the expectation of a single correct answer. But yes, my assumption that there's only a single correct answer, while reasonable in general, could be considered flawed considering the nature of the question itself.

  • source
  • parent
  • hideshow 1 child comment
  • [–] 7 points 2 months ago

    The text says “the answer” and 25% having two ways to get to it does not mean there are two answers. You can’t just say you’re avoiding assumptions and then make more assumptions that are even more outlandish.

    Ultimately, you can’t really side-step how impossible it is to answer it correctly. The answer is to write in “e) fuck off”

  • source
  • parent
  • [–] 24 points 2 months ago (2 children)

    Godel showed that self-referential problems like this are impossible to resolve. He proved that the existence of these types of problems mean that math can never completely describe reality as outlined in his incompleteness theorems -> https://en.wikipedia.org/wiki/G%C3%B6del's_incompleteness_theorems

    This is one of the drivers of post-modernist thinking. Douglas Hofstadter talks a lot about this and the nature of consciousness in his famous book Gödel, Escher, Bach -> https://en.wikipedia.org/wiki/G%C3%B6del%2C_Escher%2C_Bach

  • source
  • hideshow 2 child comments
  • [–] 5 points 2 months ago (1 child)

    as far as i understand, godel's incompleteness doesn't touch reality or ability for math to describe reality. math is very vast and we only need a fraction of it to describe reality right now. whether this description is correct is debatable but not part of godel's proof, which to me only means math can't be "summarized" into a set of "universal" rules. i may be wrong tho, godel's work is a lot to take in

  • source
  • parent
  • hideshow 1 child comment
  • [–] 3 points 2 months ago

    Yes, I described it badly and that description is inaccurate. It is technically about completeness which has a very specific definition. The implication that no formal logic system could describe even itself fully and that was a major blow to modernist “clock-maker’s universe” view of the reality.

  • source
  • parent
  • [–] 42 points 2 months ago

    Ah fuck, there goes my right to vote in Jim Crow South.

  • source
  • [–] 19 points 2 months ago (5 children)

    C) none of the possible answers is correct if picked randomly so you have 0% chance to get it right

  • source
  • hideshow 5 child comments
  • [–] 33 points 2 months ago (4 children)

    That makes 0% the correct answer, which means you had a 25% chance...

  • source
  • parent
  • hideshow 4 child comments
  • [–] 11 points 2 months ago (2 children)

    If you were to pick at random it would be 25% given there are 4 options being. However, that isn’t the same as the question about the question, which you see 25% percent twice meaning you could Be randomly right twice actually, so the answer to this question about a question is b.) 50%.

  • source
  • hideshow 2 child comments
  • [–] 10 points 2 months ago

    Around 33%?

  • source
  • [–] 8 points 2 months ago
    [–] 8 points 2 months ago

    No time to think about it! C!

  • source
  • [–] 8 points 2 months ago (1 child)

    There's a 50% chance of choosing 25%, but 25% might not be the correct answer. So I'm going 0% because it fits all of my previous life choices.

  • source
  • hideshow 1 child comment
  • [–] 3 points 2 months ago

    25% would normally be the correct answer, but since it's there twice it would be 50%, but since there's only a 1/4 chance if hitting 50%, 50% would be wrong, so it's impossible to pick the right answer, making it 0%, but since there's a 1/4 chance of selecting 0%, that makes it 25%, but 25% is there twice, so it would be 50%, but....

  • source
  • parent
  • [–] 8 points 2 months ago (1 child)

    33.3333333333333333333333333333333....

  • source
  • hideshow 1 child comment
  • [–] 7 points 2 months ago

    Classical paradox question. There is no correct answer.

  • source
  • [–] 11 points 2 months ago (1 child)

    if the correct answer is b or c, you would be right a quarter of the time

    if the correct answer is a or d you would be right half the time.

    athere are 3 correct answers. assuming the correct answer is chosen at random, you would be right 1/3x1/2 + 2/3x1/4 = 1/3 of the time

  • source
  • hideshow 1 child comment
  • [–] 5 points 2 months ago (2 children)

    So what you do is, ignore the supposed answers and and write 25% next to them.

    Problem solved!

  • source
  • hideshow 2 child comments
  • [–] 5 points 2 months ago (1 child)

    If this wasn't an extra credit question the correct answer is to drop the class.

  • source
  • hideshow 1 child comment
  • [–] 5 points 2 months ago

    The question is both self-aware and mutated. We have to burn it with fire before it learns to replicate.

  • source
  • [–] 4 points 2 months ago*

    a == d

    a + d = b

    c = burn it down

  • source
  • [–] 4 points 2 months ago
    [–] 4 points 2 months ago
    [–] 4 points 2 months ago (1 child)

    My response when I saw this again a week or two ago was "the answer is 0%, but only if you don't pick it."

  • source
  • hideshow 1 child comment
  • [–] 3 points 2 months ago (1 child)

    The answer is 50% without it being b). There is also test optimization strategy for answering F'd up questions. If 2 options are the same, they can not be intended answer of test maker.

  • source
  • hideshow 1 child comment
  • [–] 3 points 2 months ago
    [–] 3 points 2 months ago (2 children)

    I'm writing in

    e) All or none of the above.

  • source
  • hideshow 2 child comments
  • [–] 1 point 2 months ago*

    🤔... with poorly written rules it might seem proper to just make up sum shit and your color outside the lines moment might just be the one. Yet IRL rockets explode dams just burst...thermal run away yada yada soooo

  • source
  • parent
  • [–] 2 points 2 months ago (2 children)

    It can't be 25% since they appear twice, since that's disqualified it is between 50% and 0, and it isn't 0.

  • source
  • hideshow 2 child comments
  • [–] 2 points 2 months ago
    [–] 2 points 2 months ago*

    Math with no rules is a pattern. Math with rules is useful. Bouncing back and forth might refine the process?

  • source
  • [–] 2 points 2 months ago

    The struggle is real

  • source
  • [–] 2 points 2 months ago (2 children)

    A. I'll go with A. Huh, that was easy.

  • source
  • hideshow 2 child comments
  • [–] 2 points 2 months ago

    Bro. Math tests aren't real, bro.

  • source
  • [–] 2 points 2 months ago

    Quantumtative Ability

  • source
  • [–] 2 points 2 months ago
    [–] 4 points 2 months ago (7 children)

    B

    The four answers sum to 100%, so on average each answer has a 25% chance of being right, and 25% is two of the options so you have a 50% chance of picking one of those

  • source
  • hideshow 7 child comments
  • [–] 16 points 2 months ago (3 children)

    But two of those answers are the same, which reduces the number of possibilities to 3, and 33.3% isn't an option.

  • source
  • parent
  • hideshow 3 child comments
  • [–] 3 points 2 months ago (1 child)

    It doesn't reduce the numbet of possibilities. If you had 9999 true and 1 false, the chances are not 50/50

  • source
  • parent
  • hideshow 1 child comment
  • [–] 2 points 2 months ago

    25% chance that guessing an answer on a 4 answer question will be true. However, two of the answers say 25%, which is 50% of the possibilities, which now invalidates those answers from being correct. With only answer being correct, it invalidates 50% from being the right answer.

  • source
  • parent
  • [–] 7 points 2 months ago

    But if the correct answer is 50% and there is only 1/4 chance of choosing that answer at random, you then have a 25% chance of choosing it correctly. So the answer must be 25%, but there is 2/4 chance of choosing that at random so you have a 50% chance of choosing that correctly so the answer is 50%, but there is only a 1/4 chance of...

  • source
  • parent
  • [–] 5 points 2 months ago (1 child)

    What the question is asking for, is an answer where the value and the chance of choosing it at random are the same.

    None of the answers satisfy that condition, because you have a 50% chance of choosing 25%, a 25% chance of choosing 50% and a 25% chance of choosing 0%.
    None of the answers match the chance of choosing it.

  • source
  • parent
  • hideshow 1 child comment
  • [–] 2 points 2 months ago (1 child)

    C because all the answers are wrong.

  • source
  • hideshow 1 child comment
  • [–] 1 point 2 months ago

    Knowing the answers makes it impossible for random. 0%, but that’s a decision based on fact, not random. Now my brain has melted into a puddle of paradox.

  • source
  • [–] 1 point 2 months ago

    The question really is... are all questions really worth solving and what good is math if you are not solving problems. Self awarness is the pursuit of understanding the world we live in and math is a tool to do just that. Maybe the output is just misunderstood.... maybe there is no answer to some problems. The answer is 2

  • source
  • [–] 1 point 2 months ago*

    FINE! Post modernism in the chat like bla bla bla but greed is good and we were born in its shaddow. The answer is 100%

  • source
  • [–] 1 point 2 months ago

    Truth is a popularity contest created in a poorly written language. Garbage in garbage out. The answer is 666,666,666 X 10^(3000/3) or all the above

  • source
  • [–] 1 point 2 months ago

    if you were to pick randomly, the correct answer would change something you didnt pick.

    meaning the chances of a random selection picking the correct answer is zero.

    since your answer is not a random selection, you can input the answer as C.

  • source
  • [–] 1 point 2 months ago

    Mathmatical riddles. Kinda like turds that taste like skittles. News flash the people in charge don't know what they are doing.

  • source
  • [–] -1 points 2 months ago (4 children)

    Well if I pick it at random the chances of typing one of those specific char strings into the keyboard is small enough to be effectively 0%

    Creative interpretation of the question is a valid test taking method

  • source
  • hideshow 4 child comments
  • [–] 13 points 2 months ago (3 children)

    This is a paper test, though… Not sure how your keyboard factors in to that 😅

  • source
  • parent
  • hideshow 3 child comments
  • [–] -5 points 2 months ago* (last edited 2 months ago) (4 children)

    Chance == odds

    so
    
    4/4 == totalChoice_permitted 
    one_Choice = 1/4 
    
    A=25
    B=50
    C=0
    D=25
    
    if one_Choice  ==  one_Choice
    then combine 
    	reduce choices by the amount simlair plus one
    
    	A==D
    	combine
    
    3/3 == totalChoice_permitted 
    one_Choice = 1/3 
    	turn to decimal (one_Choice) x 100 == 33%
    	
    
    1 = AD =25%
    2 = B =50%
    3 = C =0%
    
    if answer not equal to (one_Choice)
    	print “the answer is not listed but the answer is in the struggle of answering a question so 0%”
    
    
    
    
    

    Not C but 0% Not all odds matter or they do. It’s a double whammy

    or do I got pie on my face?

    Is it the percentage or the place holder?

    What are the rules?

    I would think it would be safe to assume the limits are out of four choices but when two are the same.

    like linear or some shit.

    Update: I asked my local AI clown clanker and it gave this as a answer

    A clever question! Since the answer is chosen at random, each option has an equal chance of being correct. Therefore, the probability that I will be correct if I pick randomly is: 1/4 = 25% So, the correct answer is (a) 25%!

    The olgipity is very sure of A) and not D)

  • source
  • hideshow 4 child comments
  • [–] 4 points 2 months ago (2 children)

    So if you are playing Russian roulette with 5 loaded chambers, you can ignore 4 of them because the outcome is the same as one of the other chambers, meaning you have a 50/50 chance of winning. Truly astounding math.

  • source
  • parent
  • hideshow 2 child comments
  • [–] -5 points 2 months ago (2 children)

    We need to know the question do we not? OP failed here. The question could change the answer.

  • source
  • hideshow 2 child comments
  • [–] 5 points 2 months ago (1 child)

    The question is quite clear

    “What is the % chance that you can guess the right answer on this multiple choice question”

    i just wish we knew the intended answer… assuming this isn’t fake.

  • source
  • parent
  • hideshow 1 child comment
  • [–] 3 points 2 months ago

    I think it's actually 0%. Because it's impossible to answer correctly.

    But wouldn't that make the answer 25%? Which would make the answer 50%. Which would make the answer 25%. So you get stuck in an infinite loop, except that since that makes it impossible to answer, the paradox collapses, making the true answer 0%.

    Only... that would mean there's a 25% chance of guessing it. But that means it's still 0%... sorta?

    I feel like this question is more philosophical than mathematical...

  • source
  • parent